2) Atoms are Very Small The components of dry air at sea level are shown. (The oX TABLE 7.3Atmospheric Composition Near Sea Level of air at 1 atm and 25 °C, how many helium atoms would be in the tiny box? Note that 1 cm (Dry Air)* Component Mole Fraction N2 O2 Ar CO2 Ne He CH4 Kr 0.78084 0.20948 0.00934 0.000345 0.00001818 0.00000524 0.00000168 0.00000114 0.0000005 0.0000005 0.000000087 NO Xe
Answer #1

Total moles in chamber will be given by ideal gas law:

n = PV/RT

P = pressure

V = volume = 1 cm3 = 1mL = 0.001L

R = gas constant

T = temperature = 25 oC = 25 + 273 = 298K

n = 1 atm x 0.001L / 0.08206 atm.L/mol.K x 298K = 4.08933 x 10-5 moles

moles of helium = total moles x mole fraction = 4.08933 x 10-5 moles x 0.00000524 = 2.14281 x 10-10 moles

1 mole gas = 6.022 x 1023 atoms /mole

no. of He atom = 2.14281 x 10-10 mole x 6.022 x 1023 atoms /mole = 1.29 x 1014 atoms of helium